OpenStudy (anonymous):

Write the standard form of the equation of the parabola that has a vertex at (-4,11) and passes through the point (-6,15). F(x)=3-2x^4. thank you!

OpenStudy (anonymous):

what's this y=3-2x^4

OpenStudy (anonymous):

i dont know my teacher wrote it with that i got no clue what that means either :(

OpenStudy (anonymous):

It seems out of place

OpenStudy (anonymous):

i know well how do you solve it without that f(x) thing?

OpenStudy (anonymous):

y=a(x-h)^2+k y=a(x-(-4))^2+11 y=a(x+4)^2+11 Let find a Plug in a known point (-6,15) 15=a(-6+4)^2+11 15=4a+11 a=1 rewrite equation y=1(x+4)^2+11

OpenStudy (anonymous):

Standard Form Ax^2+by+c=0 Just rearrange

OpenStudy (anonymous):

oh thank you!